OpenStudy (lexy148):

a class consist of 15 members. how many commities of four can be selected?

OpenStudy (igreen):

15 / 4 = ? @lexy148

OpenStudy (anonymous):

3, because 4 cant go into 5 equally so 4x3=12 4x3=16 so three

OpenStudy (anonymous):

15*

OpenStudy (igreen):

@amburkoenigs You're not supposed to give out answers.

OpenStudy (anonymous):

oh oops

OpenStudy (perl):

15 choose 4

OpenStudy (perl):

assuming order does not count in a committee

OpenStudy (perl):

1365 different committees of 4 can be chosen from 15 people

OpenStudy (lexy148):

15/4 =3.75

OpenStudy (lexy148):

@iGreen

OpenStudy (lexy148):

how did you get 1365 @perl

OpenStudy (igreen):

Yep, you're right. 15 / 4 = 3.75. Since you can't have .75 of a person, the answer is 3.